OpenStudy (anonymous):

Brennan is making a poster for the drama club’s new production. It is a regular pentagon with side lengths of 12 inches. The school wants to put up a giant replica of the poster during athletic events. If the length of each side is 8 times the original, how many times larger is the area of the replica than the area of the original?

OpenStudy (anonymous):

@experimentX ?

OpenStudy (experimentx):

8^2 times

OpenStudy (experimentx):

if the figure is enlarged by some factor, in area it's factor squared times original in volume it's factor cubed times original

OpenStudy (anonymous):

@SMISHRA I'm confused. Could you please look at this?

OpenStudy (anonymous):

shall we go back to basic?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

 Two SIMILAR triangles here. I have enlarged it 5 times. Area A1=(1/2). bh Area A2=(1/2). 5b. 5h= 5^2 .Area of A1 Is it clear ?